5)The coordinates of the vertices of a triangle are (3,2), (9,2), (6,5).
a)Find the equations of the perpendicular bisectors of triangle.
b)Find the co ordinates of the circumcentre
1
Expert's answer
2020-10-26T19:01:32-0400

The coordinates of the vertices of a triangle are "A(3,2), B(9,2), C(6,5)."

a)

The midpoint of "AB" is "M_1\big(\frac{3+9}{2},\ \frac{2+2}{2}\big)=M_1(6,\ 2)"

The slope of "AB" is "\frac{2-2}{9-3}=\frac{0}{6}=0". It means that "AB" is a horizontal line and perpendicular bisector is a vertical line. The equation of the perpendicular bisector is "x=6" .


The midpoint of "BC" is "M_2\big(\frac{9+6}{2},\ \frac{2+5}{2}\big)=M_2(7.5,\ 3.5)" 

The slope of "BC" is "\frac{2-5}{9-6}=\frac{-3}{3}=-1". The slope of the perpendicular bisector is "m_2=-\frac{1}{-1}=1" .

The equation of the perpendicular bisector is "y-3.5=1\times (x-7.5)" or "y=x-4" .


The midpoint of "AC" is "M_3\big(\frac{3+6}{2},\ \frac{2+5}{2}\big)=M_3(4,5,\ 3.5)" 

The slope of "AC" is "\frac{2-5}{3-6}=\frac{-3}{-3}=1". The slope of the perpendicular bisector is "m_3=-\frac{1}{1}=-1" .The equation of the perpendicular bisector is "y-3.5=-1\times (x-4.5)" or "y=-x+8" .

b)

The circumcenter is the point of concurrency of perpendicular bisectors of a triangle. We need to solve system of any two bisector equations to find the coordinates of the circumcenter:

"\begin{cases}
x=6
\\
y=x-4
\end{cases}" "\begin{cases}
x=6
\\
y=2
\end{cases}"

The circumcenter is "O(6,\ 2)" .


Answers: a) "x=6" , "y=x-4" , "y=-x+8" ; b) "(6,\ 2)" .